Let z = 4(cos 135° + i sin 135°) and w = 3(cos 45° + i sin 45°). Write the rectangular form of zw.

Answer by Alan3354(69443)

Let z = 4(cos 135° + i sin 135°) and w = 3(cos 45° + i sin 45°). Write the rectangular form of zw.
---------------------
zw = 12cis(135+45) = 12cis(180)
zw = 12cos(180) = 12i*sin(180)
zw = -12